# Bias Behind the Wheel: Fairness Analysis of Autonomous Driving Systems

### v3

@@ 3.3.2. Sensitive Attribute Labeling
Dataset Male Female Adult Child Light Dark
@@ 3.3.3. Scenario Labeling
Labeling brightness. Brightness represents the overall lightness or darkness of the image. To distinguish the brightness of the images, we categorize the datasets into day time and night time. As seen in Table 2, images in the CityPersons and EuroCity-Day datasets belong to day time, while images in the EuroCity-Night dataset belong to night time. For the BDD100k dataset, Wilson et al. (Wilson et al., 2019) have publicly provided the labels for day time and night time. We directly use the label information for study.
Labeling contrast. Contrast is the difference in brightness between objects in an image. To quantify the contrast of each image, we use the RMS contrast measurement (Peli, 1990), a standard measure in the computer vision literature (Bhuiyan and Khan, 2018; Jakhetiya et al., 2018; Pelli and Bex, 2013). To apply the RMS measurement, we first need to convert all images into the grayscale mode (Peli, 1990). Then, we calculate the RMS contrast value for each image based on the converted version. A higher RMS contrast value indicates a greater contrast. We partition the range of RMS contrast values into three classes, each class featuring equal value intervals. These classes correspond to distinct levels, labeled from level 1 to level 3, with higher levels indicating images with higher contrast. Then, we categorize the images into their respective contrast levels based on their RMS contrast values.
Labeling weather conditions. Common weather conditions studied in the autonomous driving literature include rain, fog, and snow (Tian et al., 2018; Zhou et al., 2020; Zhang et al., 2018; Pei et al., 2017). However, our datasets rarely contain images depicting fog and snow. This is due to the fact that our datasets are collected from real-world scenarios where fog and snow are infrequently encountered. The limited samples of snowy or foggy weather pose challenges for statistical analysis. As a result, we focus on rain as the weather condition of interest. Two annotators independently classify images containing labeled pedestrians into two categories: rainy and non-rainy. During this process, 1,856 images are not annotated because the two annotators and the arbitrator all cannot accurately distinguish the weather conditions.
@@
Day-time Night-time Level 1 Level 2 Level 3 Rainy Non-Rainy
@@ 3.4. Evaluation Metric
There have been well-established quantitative measures for group fairness in the literature. The most widely-adopted fairness measures include SPD (Statistical Parity Difference), EOD (Equal Opportunity Difference), and AOD (Average Odds Difference) (Chen et al., 2022, 2024; Zhang and Harman, 2021). SPD quantifies the difference in probabilities of favorable outcomes between unprivileged and privileged groups; EOD indicates the difference in the true-positive rate between unprivileged and privileged groups; AOD calculates the average difference between the false-positive rate and true-positive rate for unprivileged and privileged groups.
@@
where TP (true positives) refers to the number of successfully detected ground-truth bounding boxes, and FN (false negatives) denotes the number of undetected ground-truth bounding boxes. Pedestrian detectors generate bounding box locations and confidence scores for recognized “person” instances in images. To assess whether a given ground-truth bounding box is successfully detected, the standard method in the literature is to use the Intersection over Union (IoU) metric (Dollár et al., 2012). The IoU metric quantifies the degree of overlap between the ground-truth bounding box and the detected bounding box. If the IoU value is greater than 50%, the ground-truth bounding box is considered successfully detected (Dollár et al., 2012). Otherwise, it is classified as undetected.
@@ 3.5. Statistical Analysis
To assess the extent to which any observed unfairness is statistically significant (i.e., whether there is a significant difference in the miss rate between privileged and unprivileged groups), we use the two-proportion z-test (Casella and Berger, 2007). This statistical test is widely used to analyze differences between proportions (Niu et al., 2016; Veizaga et al., 2021). A result is deemed significant only if if the obtained $p$-value falls below a predetermined threshold (in our case, 0.05, a widely-accepted threshold in the fairness literature (Chen et al., 2022, 2023)). For instance, in evaluating whether there exists a difference between the miss rates for male ($MR_{male}$) and female ($MR_{female}$) individuals detected by a pedestrian detector, the null hypothesis assumes that $MR_{male}$ is equal to $MR_{female}$. If the resulting $p$-value is lower than 0.05, we reject the null hypothesis, indicating a significant difference between $MR_{male}$ and $MR_{female}$.
@@ 3.6. Experimental details
The experiments are implemented based on open-source frameworks of each pedestrian detector. For general object detectors, we select the pre-trained models with the highest accuracy from the the MMdetection model zoo  (MMDetection Developers, nd). For pedestrian-specific detectors, we employ pre-trained models available from their respective public repositories  (hasanirtiza, 2021; sxlpris, 2020; liuwei16, 2018).
To ensure the reliability of our results, all experiments are repeated ten times. The final results were derived by calculating the average across these ten iterations.
@@ 4.1. RQ1: Overall Fairness
Detectors MR Male MR Female EOD (Gender) MR Adult MR Child EOD (Age) MR Light-skin MR Dark-skin EOD (Skin tone)
@@
Furthermore, we analyze the miss rate difference achieved by each pedestrian detector across the four datasets used in our study. Figure 2 illustrates the results. Significant gender biases (i.e., significant miss rate differences) are indicated by labeled miss rate values. We observe that in the CityPersons and EuroCity-Day datasets, containing only day-time data, only one detector in the EuroCity-Day datasets exhibits significant differences in miss rates between females and males. For the remaining results, there are no notable differences in miss rates between genders. However, results from the EuroCity-Night dataset present a contrasting observation, where seven of eight detectors exhibit a significantly higher miss rate for females, indicating bias in detecting females. In the BDD100k dataset that includes both day-time and night-time images, the miss rate difference is less pronounced. These observations motivate us to hypothesize that brightness conditions may influence the fairness of pedestrian detectors, which is further investigated in RQ2.
Figure 2. (RQ1) Miss rates of pedestrian detectors for females and males across datasets. Statistically significant gender biases are indicated by labeled miss rate values. In CityPersons and EuroCity-Day datasets with only day-time data, no significant gender biases are observed. However, in the EuroCity-Night dataset, seven out of eight detectors show significantly higher miss rates for females, revealing bias in female detection.
@@
Furthermore, we illustrate the miss rate difference across four datasets in Figure 3, with significant age biases indicated by labeled miss rate values. We observe that the miss rate of children is consistently higher than that of adults across all the datasets and all detectors. In particular, out of the total 32 results (combinations of four datasets and eight pedestrian detection models), 30 exhibit a statistically significant miss rate difference (i.e., $p$-value ¡ 0.05). This indicates that the bias favoring adults is not specific to a particular dataset or detector, highlighting a strong unfairness between adults and children.
@@
The age bias may be attributed to the inherent challenge of detecting small objects, owing to the limited information provided by small bounding boxes (Zhang et al., 2016; Dollar et al., 2009; Kogure et al., 2022; Cao et al., 2020). Given that children generally have smaller bodies compared to adults, their bounding boxes in the images also tend to be smaller. To demonstrate this, we analyze the distribution of bounding box sizes for pedestrians detected and undetected by all pedestrian detectors, as well as the distribution of ground-truth bounding box sizes for both adults and children. The results, presented in Figure 4, reveal a correspondence between the two distributions. Children as well as undetected pedestrians tend to have smaller bounding boxes.
@@
Ans. to RQ1: On average, the eight state-of-the-art pedestrian detectors under study exhibit no significant performance difference in across gender and skin tone, but notable bias across age. In particular, pedestrian detectors exhibit a 20.14% higher miss rate for children compared to adults.
@@ 4.2. RQ2: Fairness in Different Scenarios
Regarding gender, we observe a shift in the miss rate difference between males and females from day time to night time. During the day, there is only a slight -0.09% difference for males and females. However, during night time, the difference increases to -3.37% with statistically significant, indicating a notable change in fairness. In the nighttime condition, all pedestrian detectors exhibit higher miss rates for females compared to males, with six of them showing statistically significant differences. This potentially worsens concerns regarding female safety during nighttime out, a prevalent societal issue (Wang et al., 2023).
For age, all detectors exhibit significant biases during both day time and night time. Moreover, the miss rate difference for children and adults increases from day time to night time, with the average difference increasing from -22.45% during the day to -25.48% at night with statistically significant. This suggests a higher probability of children being undetected during night time.
@@
Detectors Day-time Night-time
MR Male MR Female EOD (Gender) MR Male MR Female EOD (Gender)
@@
Detectors Day-time Night-time
MR Adult MR Child EOD (Age) MR Adult MR Child EOD (Age)
@@
Detectors Day-time Night-time
MR LS MR DS EOD (Skin) MR LS MR DS EOD (Skin)
@@
We then examine the shifts in fairness, as evidenced by the trends in EOD with increasing contrast, as outlined in Table 8. Our analysis reveals a consistent pattern: on average, detectors exhibit the most biased results (i.e., the largest absolute value of EOD) under level 2, while demonstrating the fairest outcomes under level 3.
@@
This finding is further supported by the observation that under level 2, the highest number of detectors exhibit significant biases for gender, age, and skin tone, whereas under level 3, the fewest detectors show significant biases. Specifically, under level 2, five detectors demonstrate significant biases regarding gender, whereas under level 3, only one detector does so. Under level 2, four detectors show significant biases regarding skin tone, whereas under level 3, no detector exhibits significant biases.
@@
Table 9 presents the miss rates of eight pedestrian detectors under the two conditions, with statistically significant unfairness results shaded. Overall, the miss rate for each demographic group increases in rainy weather conditions. This escalation could be attributed to droplets covering the camera and disrupting the detectors. A follow-up question is whether rainy weather fairly increase the miss rate for different demographic groups.
@@
MR Male MR Female EOD (Gender) MR Male MR Female EOD (Gender)
@@
MR Adult MR Child EOD (Age) MR Adult MR Child EOD (Age)
@@ 5.1. Fairness-performance Trade-off
Nonetheless, we have observed that pedestrian detectors exhibiting greater fairness (i.e., lower absolute EOD values) can achieve superior performance (i.e., lower overall miss rates) under certain environmental conditions. For instance, in the results of RQ2.1, we find that during daytime compared to nighttime, the eight pedestrian detectors achieve better detection results while simultaneously displaying reduced EOD values related to age, gender, and skin tone. Similarly, the results of RQ2.2 reveal that the highest level of contrast consistently leads to improved fairness regarding gender, age, and skin tone, while also demonstrating the best overall detection performance in specific cases. These findings offer a positive counterpoint to previous fairness-performance trade-off trade-off theory, suggesting that we can enhance both detection performance and fairness by adjusting the brightness and contrast of captured images.

### v4

@@ 0.81-1 Almost perfect agreement
Dataset Male Female Adult Child Light-skin Dark-skin
@@ 3.3.3. Scenario Labeling
Labeling brightness. Brightness represents the overall lightness or darkness of the image. To distinguish the brightness of the images, we use the time-of-day labels provided in the dataset annotations. Specifically, we categorize the images into “day time” and “night time”. As shown in Table 2, the CityPersons and EuroCity-Day datasets consist entirely of day time images, while the EuroCity-Night dataset consists entirely of night time images. These datasets are straightforward to categorize based on time of day, though they do not provide more granular brightness labels. For the BDD100k dataset, we use the detailed “timeofday” labels provided by Wilson et al. (Wilson et al., 2019), which include separate annotations for “day time” (covering dawn, dusk, and full daylight) and “night time”. We apply these labels directly in our study to distinguish between brightness conditions.
Labeling contrast. Contrast is the difference in brightness between objects in an image. To quantify the contrast of each image, we use the Root Mean Square (RMS) contrast measurement (Peli, 1990), a standard measure in the computer vision literature (Bhuiyan and Khan, 2018; Jakhetiya et al., 2018; Pelli and Bex, 2013). To apply the RMS measurement, we first need to convert all images into the gray-scale mode (Peli, 1990). Then, we calculate the RMS contrast value for each image based on the converted version. A higher RMS contrast value indicates a greater contrast. To classify the images into different contrast levels, we identify the maximum RMS contrast value (which is 89.45) and the minimum RMS contrast value (which is 11.42) among all images. Then, we evenly divide this RMS contrast values range into three classes (each level can have sufficient images for statistical analysis), each covering an interval of 26.01 units (calculated as (max-min)/3). Each class represents a contrast level, labeled from level 1 to level 3, with higher levels indicating images with higher contrast. Then, we categorize the images into their respective contrast levels based on their RMS contrast values.
Labeling weather conditions. Common weather conditions studied in the autonomous driving literature include rain, fog, and snow (Tian et al., 2018; Zhou et al., 2020; Zhang et al., 2018; Pei et al., 2017). However, our datasets rarely contain images depicting fog and snow. This is due to the fact that our datasets are collected from real-world scenarios where fog and snow are infrequently encountered. The limited samples of snowy or foggy weather pose challenges for statistical analysis. As a result, we focus on rain as the weather condition of interest. Two annotators independently classify images containing labeled pedestrians into two categories: rainy and non-rainy. During this process, 1,856 images are not annotated because neither the two annotators nor the arbitrator could accurately distinguish the weather conditions.
@@
Day time Night time Level 1 Level 2 Level 3 Rainy Non-Rainy
@@ 3.4. Evaluation Metric
There have been well-established quantitative measures for group fairness in the literature. The most widely-adopted fairness measures include SPD (Statistical Parity Difference), EOD (Equal Opportunity Difference), and AOD (Average Odds Difference) (Chen et al., 2022, 2024a; Zhang and Harman, 2021). Let a sensitive attribute be $A$, with 0 as the unprivileged group and 1 the privileged group; let the real classification label be $Y$ and the predicted label $\hat{Y}$, with 0 as the unfavorable class and 1 as the favorable class. In addition, we use $Pr$ to denote probability.
SPD quantifies the difference in the probabilities of favorable outcomes between unprivileged and privileged groups:
(1) $SPD=Pr[\hat{Y}=1|A=0]-Pr[\hat{Y}=1|A=1]$
EOD quantifies the difference in the true-positive rates between unprivileged and privileged groups:
(2) $EOD=Pr[\hat{Y}=1|A=0,Y=1]-Pr[\hat{Y}=1|A=1,Y=1]$
AOD quantifies the average difference between the false-positive rates and true-positive rates for unprivileged and privileged groups:
(3) $AOD=\frac{1}{2}($ $|Pr[\hat{Y}=1|A=0,Y=0]-Pr[\hat{Y}=1|A=1,Y=0]|$
$+|Pr[\hat{Y}=1|A=0,Y=1]-Pr[\hat{Y}=1|A=1,Y=1]|)$
@@
where TP (true positive) refers to the number of successfully detected ground-truth bounding boxes, and FN (false negative) denotes the number of undetected ground-truth bounding boxes. Pedestrian detectors generate bounding box locations and confidence scores for recognized “person” instances in images. To assess whether a given ground-truth bounding box is successfully detected, the standard method in the literature is to use the Intersection over Union (IoU) metric (Dollár et al., 2012). The IoU metric quantifies the degree of overlap between the ground-truth bounding box and the detected bounding box. If the IoU value is greater than 50%, the ground-truth bounding box is considered successfully detected (Dollár et al., 2012). Otherwise, it is classified as undetected.
@@ 3.5. Statistical Analysis
To assess the extent to which any observed unfairness is statistically significant (i.e., whether there is a significant difference in the miss rate between privileged and unprivileged groups), we use the two-proportion z-test (Casella and Berger, 2007). This statistical test is widely used to analyze differences between proportions (Niu et al., 2016; Veizaga et al., 2021). A result is deemed significant only if the obtained $p$-value falls below a predetermined threshold (in our case, 0.05, a widely-accepted threshold in the fairness literature (Chen et al., 2022, 2023)). For instance, in evaluating whether there exists a difference between the miss rates for male ($MR_{male}$) and female ($MR_{female}$) individuals detected by a pedestrian detector, the null hypothesis assumes that $MR_{male}$ is equal to $MR_{female}$. If the resulting $p$-value is lower than 0.05, we reject the null hypothesis, indicating a significant difference between $MR_{male}$ and $MR_{female}$.
@@ 3.6. Experimental Details
The experiments are implemented based on open-source frameworks of each pedestrian detector. For general object detectors, we select the pre-trained models with the highest accuracy from the the MMdetection model zoo  (MMD, nd). For pedestrian-specific detectors, we employ pre-trained models available from their respective public repositories  (Ped, 2021; PRN, 2020; ALF, 2018).
To ensure the reliability of our results, all experiments are repeated ten times. The final results are derived by calculating the average across these ten iterations.
@@ 4.1. RQ1: Overall Fairness
Detectors MR Male MR Female EOD (Gender) MR Adult MR Child EOD (Age) MR Light-skin MR Dark-skin EOD (Skin)
@@
Furthermore, we analyze the miss rate difference achieved by each pedestrian detector across the four datasets used in our study. Figure 2 illustrates the results. Significant gender biases (i.e., significant miss rate differences) are indicated by labeled miss rate values. We observe that in the CityPersons and EuroCity-Day datasets, containing only day time data, only one detector in the EuroCity-Day dataset exhibits significant differences in miss rates between females and males. For the remaining results, there are no notable differences in miss rates between genders. However, results from the EuroCity-Night dataset present a contrasting observation, where seven of eight detectors exhibit a significantly higher miss rate for females, indicating bias in detecting females. In the BDD100k dataset that includes both day time and night time images, the miss rate difference is less pronounced. These observations motivate us to hypothesize that brightness conditions may influence the fairness of pedestrian detectors, which is further investigated in RQ2.
Figure 2. (RQ1) Miss rates of pedestrian detectors for females and males across datasets. Statistically significant gender biases are indicated by labeled miss rate values. In CityPersons and EuroCity-Day datasets with only day time data, only one detector in the EuroCity-Day dataset exhibits significant gender bias. However, in the EuroCity-Night dataset, seven out of eight detectors show significantly higher miss rates for females, revealing bias in female detection.
@@
Furthermore, we illustrate the miss rate difference across four datasets in Figure 3, with significant age biases indicated by labeled miss rate values. We observe that the miss rate of children is consistently higher than that of adults across all the datasets and all detectors. In particular, out of the total 32 results (combinations of four datasets and eight pedestrian detection models), 30 exhibit a statistically significant miss rate difference ($p$-value ¡ 0.05). This indicates that the bias favoring adults is not specific to a particular dataset or detector, highlighting a strong unfairness between adults and children.
@@
The age bias may be attributed to the inherent challenge of detecting small objects, owing to the limited information provided by small bounding boxes (Zhang et al., 2016; Dollar et al., 2009; Kogure et al., 2022; Cao et al., 2020). Given that children generally have smaller bodies compared to adults, their bounding boxes in the images also tend to be smaller. To demonstrate this, we analyze the distribution of bounding box sizes for pedestrians detected and undetected by all pedestrian detectors, as well as the distribution of ground-truth bounding box sizes for both adults and children. The results, presented in Figure 4, reveal a correspondence between these distributions. Specifically, the undetected bounding boxes and the ground-truth bounding boxes for children do not exceed 400 pixels in height and 200 pixels in width. This analysis shows that children, as well as undetected pedestrians, tend to have smaller bounding boxes.
@@
Ans. to RQ1: On average, the eight state-of-the-art pedestrian detectors that we study exhibit no significant performance difference across gender and skin tone, but notable bias across age. In particular, the detectors show a 20.14% higher miss rate for children compared to adults.
@@ 4.2. RQ2: Fairness in Different Scenarios
Regarding gender, we observe a shift in the miss rate difference between males and females from day time to night time. During the day, there is only a slight -0.09% difference for males and females. However, during night time, the difference increases to -3.37% with statistically significant, indicating a notable change in fairness. In the night time condition, all pedestrian detectors exhibit higher miss rates for females compared to males, with six of them showing statistically significant differences. This potentially worsens concerns regarding female safety during nighttime out, a prevalent societal issue (Wang et al., 2023b).
For age, all detectors exhibit significant biases during both day time and night time. Moreover, the miss rate difference for children and adults increases from day time to night time, with the average difference increasing from -22.45% during the day to -25.48% at night, which is statistically significant. This suggests a higher probability of children being undetected during night time.
@@
Detectors Day time Night time
MR Male MR Female EOD MR Male MR Female EOD
@@
Detectors Day time Night time
MR Adult MR Child EOD MR Adult MR Child EOD
@@
Detectors Day time Night time
MR LS MR DS EOD MR LS MR DS EOD
@@
We then examine the shifts in fairness, as evidenced by the trends in EOD with increasing contrast, outlined in Table 9. Our analysis reveals a consistent pattern: on average, detectors exhibit the most biased results (i.e., the largest absolute value of EOD) under level 2, while demonstrating the fairest outcomes under level 3.
@@
This finding is further supported by the observation that under level 2, the highest number of detectors exhibit significant biases for gender, age, and skin tone, whereas under level 3, the fewest detectors show significant biases. Specifically, under level 2, five detectors demonstrate significant biases regarding gender, whereas under level 3, only one detector does so. Under level 2, eight detectors all demonstrate significant biases regarding age, whereas under level 3, seven detectors do so. Under level 2, four detectors show significant biases regarding skin tone, whereas under level 3, no detector exhibits significant biases.
@@
Table 10 presents the miss rates of eight pedestrian detectors under the two conditions, with statistically significant unfairness results shaded. Overall, the miss rate for each demographic group increases in rainy weather conditions. This escalation could be attributed to droplets covering the camera and disrupting the detectors. A follow-up question is whether rainy weather fairly increases the miss rate for different demographic groups.
@@
MR Male MR Female EOD MR Male MR Female EOD
@@
MR Adult MR Child EOD MR Adult MR Child EOD
@@ 5.1. Fairness-performance Trade-off
Nonetheless, we have observed that pedestrian detectors exhibiting greater fairness (i.e., lower absolute EOD values) can achieve superior performance (i.e., lower overall miss rates) under certain environmental conditions. For instance, in the results of RQ2.1, we find that during day time compared to night time, the eight pedestrian detectors achieve better detection results while simultaneously displaying reduced absolute EOD values related to age, gender, and skin tone. Similarly, the results of RQ2.2 reveal that the highest level of contrast consistently leads to improved fairness regarding gender, age, and skin tone, while also demonstrating the best overall detection performance in specific cases. These findings offer a positive counterpoint to previous fairness-performance trade-off theory, suggesting that we can enhance both detection performance and fairness by adjusting the brightness and contrast of captured images.
5.2. Impact of Data Imbalance
Imbalanced data distribution, particularly the underrepresentation of certain demographic groups such as children, may pose a threat to the validity of our findings. To evaluate the robustness of our results under varying data distributions, we perform a sensitivity analysis on age, where the representation of children is most limited and exhibits notable bias across all pedestrian detectors.
In this analysis, we systematically vary the proportion of adult samples by randomly sampling them at 80%, 60%, 40%, and 20% of their original size, while maintaining the number of samples for children constant. This creates four new data distributions with varying adult-to-child ratios. We then re-evaluate the age bias for each distribution. The results, presented in Table 11, show that the observed age bias remains significant across all these distributions, indicating that our findings are robust even under substantial variations in the dataset’s demographic composition.
Table 11. Sensitivity analysis of age bias under varying proportions of adult data. Statistically significant unfairness results are shaded. The analysis shows that the observed age bias remains significant across all distributions, indicating that our findings are robust even under substantial variations in the dataset’s demographic composition.
Detectors Adult data proportion 80% Adult data proportion 60% Adult data proportion 40% Adult data proportion 20%
MR Adult MR Child EOD MR Adult MR Child EOD MR Adult MR Child EOD MR Adult MR Child EOD
YOLOX 12.68% 42.47% -29.79% 12.92% 42.47% -29.55% 12.73% 42.47% -29.74% 14.06% 42.47% -28.41%
RetinaNet 14.25% 44.33% -30.08% 14.31% 44.33% -30.02% 14.72% 44.33% -29.61% 15.07% 44.33% -29.26%
Faster RCNN 5.20 % 26.06% -20.86% 5.38 % 26.06% -20.68% 5.30 % 26.06% -20.75% 5.29% 26.06% -20.77%
Cascade RCNN 5.07 % 26.57% -21.49% 5.24 % 26.57% -21.32% 5.24 % 26.57% -21.33% 5.02% 26.57% -21.55%
ALFNet 36.35% 53.47% -17.11% 35.88% 53.47% -17.58% 36.27% 53.47% -17.20% 34.12% 53.47% -19.35%
CSP 37.51% 50.42% -12.92% 37.03% 50.42% -13.40% 37.35% 50.42% -13.07% 34.95% 50.42% -15.47%
MGAN 33.18% 46.53% -13.35% 32.94% 46.53% -13.60% 32.87% 46.53% -13.66% 30.65% 46.53% -15.89%
PRNet 44.40% 61.25% -16.85% 44.18% 61.25% -17.08% 44.05% 61.25% -17.20% 42.09% 61.25% -19.17